# GM-PRM: A Generative Multimodal Process Reward Model for Multimodal Mathematical Reasoning

### v1

@@ 2 Related Work
After achieving certain results and progress in the research of language modality in PRMs, research on PRMs has also begun to shift towards multimodal tasks. M-STAR (Liu et al. 2024a) proposes and implements multimodal PRM on multimodal problems. URSA (Luo et al. 2025) constructs a dataset by inserting errors and utilizes it to train a multimodal PRM. VisualPRM (Wang et al. 2025b) not only uses MC estimation to construct a multimodal VisualPRM400K training dataset, but also proposes a benchmark for multimodal PRMs called VisualProcessBench, which is entirely annotated by humans. Athena-PRM (Wang et al. 2025a) proposes using prediction consistency between strong and weak completers to enhance the quality of automatically annotated data based on MC estimation and improving multimodal PRM by ORM initialization and negative data up-sampling. Moreover, PRM-BAS (Hu et al. 2025), $M$-PRM (Du et al. 2025) and DreamPRM (Cao et al. 2025) also improve the capability of multimodal PRMs. Although several studies have explored multimodal PRMs, applying them to multimodal tasks effectively remains certain challenges, such as insufficient interpretability of the labels assigned to each reasoning step and the inability to correct identified erroneous steps. In our work, we introduce a generative multimodal PRM, GM-PRM to solve the above problems.
@@ 3 Methodology
In this section, we first describe how to utilize PRMs and generative PRMs combined with the BoN method to improve the performance of policy models for mathematical problems in Section 3.1. Then, we introduce our process to implement multimodal generative PRM, including data construction and model training in Section 3.2. Finally, we propose a novel Refined-BoN framework for PRMs to enhance its performance beyond traditional BoN method in Section $3.$
@@ 3.1 PRMs for Mathematical Problem
where $f_{\text{{GM-PRM}}}:(\cdot)$ represents the inference of GM-PRM, II denotes the image of the mathematical problem, $c_{i}$ denotes the critique of the $i$-th reasoning step $r_{i}$, and $j_{i}$ denotes the textual judgment assigned to the $i$-th reasoning step $r_{i}$.
@@ 3.2 Data Construction
where RSRS denotes refined step of the first error step in a reasoning process.
@@ 3.3 Refined-BoN Process
After applying the Refined-BoN process, we obtain NN solutions for each problem, each accompanied by step-level correctness judgments. Moreover, we divide all the solutions into two categories: one where GM-PRM judges that it contains incorrect steps, and the other where GM-PRM judges that all its steps are correct. Furthermore, we take the probability of GM-PRM generating the associated “Correct” and “Incorrect” tokens as the score of each step.
Among the NN generated solutions, if there exist solutions in which all reasoning steps are judged correct, we calculate the average of the scores of all steps in these solutions as the overall score of the solution, and select the solution with the highest average score as the optimal solution.
For N solutions to a problem, if GM-PRM determines that all NN solutions contain incorrect steps, we calculate the average score of all steps in a solution as the overall score of the solution, and select the solution with the highest overall score as the final answer.
@@ 4.3 Hyperparameter & Ablation Study
Figure 4 depicts WeMath accuracy as the number of sampled solutions per problem (NN) increases from 1 to 8. Across all three backbones—MiniCPM‑V‑2.6‑8B, Qwen2.5‑VL‑7B, and InternVL3‑78B—both GM-PRM and the Self‑Consistency (SC) baseline benefit from a larger sampling budget, yet GM-PRM exhibits a noticeably steeper growth curve.

### v2

@@ 2 Related Work
After achieving certain results and progress in the research of language modality in PRMs, research on PRMs has also begun to shift towards multimodal tasks. M-STAR (Liu et al. 2024a) proposes and implements multimodal PRM on multimodal problems. URSA (Luo et al. 2025) constructs a dataset by inserting errors and utilizes it to train a multimodal PRM. VisualPRM (Wang et al. 2025b) not only uses MC estimation to construct a multimodal VisualPRM400K training dataset, but also proposes a benchmark for multimodal PRMs called VisualProcessBench, which is entirely annotated by humans. Athena-PRM (Wang et al. 2025a) proposes using prediction consistency between strong and weak completers to enhance the quality of automatically annotated data based on MC estimation and improving multimodal PRM by ORM initialization and negative data up-sampling. Moreover, PRM-BAS (Hu et al. 2025), MM-PRM (Du et al. 2025) and DreamPRM (Cao et al. 2025) also improve the capability of multimodal PRMs. Although several studies have explored multimodal PRMs, applying them to multimodal tasks effectively remains certain challenges, such as insufficient interpretability of the labels assigned to each reasoning step and the inability to correct identified erroneous steps. In our work, we introduce a generative multimodal PRM, GM-PRM to solve the above problems.
@@ 3 Methodology
In this section, we first describe how to utilize PRMs and generative PRMs combined with the BoN method to improve the performance of policy models for mathematical problems in Section 3.1. Then, we introduce our process to implement multimodal generative PRM, including data construction and model training in Section 3.2. Finally, we propose a novel Refined-BoN framework for PRMs to enhance its performance beyond traditional BoN method in Section 3.3.
@@ 3.1 PRMs for Mathematical Problem
where $f_{\text{{GM-PRM}}}:(\cdot)$ represents the inference of GM-PRM, $I$ denotes the image of the mathematical problem, $c_{i}$ denotes the critique of the $i$-th reasoning step $r_{i}$, and $j_{i}$ denotes the textual judgment assigned to the $i$-th reasoning step $r_{i}$.
@@ 3.2 Data Construction
where $RS$ denotes refined step of the first error step in a reasoning process.
@@ 3.3 Refined-BoN Process
After applying the Refined-BoN process, we obtain $N$ solutions for each problem, each accompanied by step-level correctness judgments. Moreover, we divide all the solutions into two categories: one where GM-PRM judges that it contains incorrect steps, and the other where GM-PRM judges that all its steps are correct. Furthermore, we take the probability of GM-PRM generating the associated “Correct” and “Incorrect” tokens as the score of each step.
Among the $N$ generated solutions, if there exist solutions in which all reasoning steps are judged correct, we calculate the average of the scores of all steps in these solutions as the overall score of the solution, and select the solution with the highest average score as the optimal solution.
For N solutions to a problem, if GM-PRM determines that all $N$ solutions contain incorrect steps, we calculate the average score of all steps in a solution as the overall score of the solution, and select the solution with the highest overall score as the final answer.
@@ 4.3 Hyperparameter & Ablation Study
Figure 4 depicts WeMath accuracy as the number of sampled solutions per problem ($N$) increases from 1 to 8. Across all three backbones—MiniCPM‑V‑2.6‑8B, Qwen2.5‑VL‑7B, and InternVL3‑78B—both GM-PRM and the Self‑Consistency (SC) baseline benefit from a larger sampling budget, yet GM-PRM exhibits a noticeably steeper growth curve.